# Efficiently Seeking Flat Minima for Better Generalization in Fine-Tuning Large Language Models and Beyond

### v2

@@ A.2 Proof of Theorem 1
The update process of the Flat-LoRA is as follows:
${\tilde{\mathbf{x}}_{t}}={{\mathbf{x}}_{t}}+\rho\frac{1}{s}\frac{{{\mathbf{G}_{t}}}}{{\left\|{{\mathbf{G}_{t}}}\right\|}}{\mathbf{y}_{t}}^{+}$ $,\quad{\tilde{\mathbf{y}}_{t}}={{\mathbf{y}}_{t}}$ (27)
@@
$={\mathbf{x}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})-({\mathbf{y}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}^{\top}({{\mathbf{x}}_{t}}+\rho\frac{1}{s}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{\mathbf{G}_{t}}\right\|}_{F}}}}\mathbf{y}_{t}^{+})))$
$={\mathbf{x}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})-({\mathbf{y}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{{x}}_{t}}+\rho\frac{1}{s}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}))$
$={\mathbf{x}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}}-({\mathbf{x}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}}){{}^{\top}}-\rho\frac{1}{s}{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}$
$=-\rho\frac{1}{s}{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}$
$=-\rho\frac{1}{s}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{G}_{{t}}}\mathbf{y}_{t}^{+}}\right]$
@@
$=-\rho\frac{1}{s}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{G}_{{t}}}\mathbf{y}_{t}^{+}}\right]$
$=-\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})}^{\top}}\mathbf{{g}_{x}}{{(\mathbf{y}_{t}^{\top})}^{+}}\mathbf{y}_{t}^{+}}\right]$
$=-\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{{}\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{(\mathbf{y}_{t}^{+})}^{\top}}\mathbf{y}_{t}^{+}}\right]$
$=-\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
$=-\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{g}_{{{\mathbf{{\tilde{x}}}}_{t}}}}({\mathbf{\tilde{y}}_{t}}^{+})^{\top}{\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
$=-\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
@@
$=\left|{-\rho\frac{1}{s}\frac{{{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}}{{{{\left\|{{\mathbf{{g}}_{\mathbf{x}}}{{(\mathbf{y}_{t}^{+})}^{\top}}}\right\|}_{F}}}}\left[{{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}}\right]}\right|$
$=\left|{-\rho\frac{1}{s}\frac{{{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}}{{\left\|{\mathbf{{g}_{x}}}\right\|\left\|{{{(\mathbf{y}_{t}^{+})}^{\top}}}\right\|}}\left[{{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}}\right]}\right|$
$=\left|{\rho\frac{1}{s}\frac{{\left\|{\mathbf{y}_{t}^{+}}\right\|}}{{\left\|{\mathbf{{g}_{x}}}\right\|}}\left[{{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}}\right]}\right|$
@@ B.1 Details on datasets
$\rho$ for EFlat-LoRA 0.6
$\beta$ for EFlat-LoRA 0.99
@@
$\rho$ for Flat-LoRA 0.3
$\beta$ for Flat-LoRA 0.95
$\rho$ for EFlat-LoRA 0.3
$\beta$ for EFlat-LoRA 0.95
@@
$\rho$ for Flat-LoRA 0.1
$\beta$ for Flat-LoRA 0.99
$\rho$ for EFlat-LoRA 0.1
$\beta$ for EFlat-LoRA 0.99
@@ B.2 Details on models
$\rho$ for Flat-LoRA 0.6
$\beta$ for Flat-LoRA 0.99
$\rho$ for EFlat-LoRA 0.1, 0.2, 0.5
$\beta$ for EFlat-LoRA 0.99
@@
Algorithm 1 Pseudocode of the Flat-LoRA
@@ B.3 Details on hyperparameters
$\rho$ for EFlat-LoRA 0.2,0.6,1
$\beta$ for EFlat-LoRA 0.99
@@
Figure 4: Evolution of balancedness across layers during training with Adam and Flat-LoRA.
@@ C. Algorithm
The two algorithms presented in 1 and 2 describe the training procedures of the proposed Flat-LoRA and its accelerated variant EFlat-LoRA.
Algorithm 2 Pseudocode of the EFlat-LoRA
@@ D.1 The approximate ability of EMA perturbation
D.2 The change in balancedness during Flat-LoRA training
We consider few shot learning with LoRA on RoBERTa-large. For dataset MNLI, 1st, 12th and 24th query layers’ $2|B_{t,l}|$ are plotted, where $t$ denotes the iteration and $l$ denotes the layer index. The layers are chosen to represent early, middle, and final stages of RoBERTa. Balancedness of Flat-LoRA and Adam on different layers are plotted in Fig. 4. Balancedness may increase or decrease across different layers. As shown in Fig. 4, the balancedness of Flat-LoRA in the first query layer of RoBERTa-large gradually decreases during training, while in the 12th layer, it first decreases and then increases. In contrast, the balancedness in the 24th layer continuously increases. An increase typically occurs when parameter magnitudes in both low-rank subspaces grow simultaneously. This behavior can be influenced by factors such as the learning rate, optimization algorithm, weight decay, and other regularization strategies. Despite these occasional increases, Flat-LoRA generally maintains lower balancedness than Adam in most layers, suggesting its capacity to induce implicit regularization during training.

### v3

@@ A.2 Proof of Theorem 1
The update process of the FMLoRA is as follows:
${\tilde{\mathbf{x}}_{t}}={{\mathbf{x}}_{t}}+\rho\frac{1}{s}\frac{{{\mathbf{G}_{t}}}}{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}{\mathbf{y}_{t}}^{+}$ $,\quad{\tilde{\mathbf{y}}_{t}}={{\mathbf{y}}_{t}}$ (27)
@@
$=-{\mathbf{x}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})+({\mathbf{y}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}^{\top}({{\mathbf{x}}_{t}}+\rho\frac{1}{s}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{\mathbf{G}_{t}}\right\|}_{F}}}}\mathbf{y}_{t}^{+})))$
$=-{\mathbf{x}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})+({\mathbf{y}_{t}}^{\top}({\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{{x}}_{t}}+\rho\frac{1}{s}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}))$
$=-{\mathbf{x}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}}+({\mathbf{x}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}}){{}^{\top}}+\rho\frac{1}{s}{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}$
$=\rho\frac{1}{s}{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}\frac{{{\mathbf{G}_{{t}}}}}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\mathbf{y}_{t}^{+}$
$=\rho\frac{1}{s}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{G}_{{t}}}\mathbf{y}_{t}^{+}}\right]$
@@
$=\rho\frac{1}{s}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{y}_{t}}^{\top}{\mathbf{\tilde{G}}_{t}}^{\top}{\mathbf{G}_{{t}}}\mathbf{y}_{t}^{+}}\right]$
$=\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})}^{\top}}\mathbf{{g}_{x}}{{(\mathbf{y}_{t}^{\top})}^{+}}\mathbf{y}_{t}^{+}}\right]$
$=\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{{}\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{(\mathbf{y}_{t}^{+})}^{\top}}\mathbf{y}_{t}^{+}}\right]$
$=\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{\tilde{G}}_{t}}{\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
$=\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{{({\mathbf{g}_{{{\mathbf{{\tilde{x}}}}_{t}}}}({\mathbf{\tilde{y}}_{t}}^{+})^{\top}{\mathbf{y}_{t}})}^{\top}}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
$=\rho\frac{1}{{{s^{2}}}}\frac{1}{{{{\left\|{{\mathbf{G}_{t}}}\right\|}_{F}}}}\left[{{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}\right]$
@@
$=\left|{\rho\frac{1}{s}\frac{{{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}}{{{{\left\|{{\mathbf{{g}}_{\mathbf{x}}}{{(\mathbf{y}_{t}^{+})}^{\top}}}\right\|}_{F}}}}({{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}})}\right|$
$=\left|{\rho\frac{1}{s}\frac{{{{\left\|{\mathbf{y}_{t}^{+}}\right\|}^{2}}}}{{\left\|{\mathbf{{g}_{x}}}\right\|\left\|{{{(\mathbf{y}_{t}^{+})}^{\top}}}\right\|}}({{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}})}\right|$
$=\left|{\rho\frac{1}{s}\frac{{\left\|{\mathbf{y}_{t}^{+}}\right\|}}{{\left\|{\mathbf{{g}_{x}}}\right\|}}({{\mathbf{g}_{{{{\mathbf{\tilde{x}}}}_{t}}}}^{\top}\mathbf{g_{x}}})}\right|$
@@ B.1 Details on datasets
$\rho$ for EFMLoRA 0.6
$\beta$ for EFMLoRA 0.99
@@
$\rho$ for FMLoRA 0.3
$\beta$ for FMLoRA 0.95
$\rho$ for EFMLoRA 0.3
$\beta$ for EFMLoRA 0.95
@@
$\rho$ for FMLoRA 0.1
$\beta$ for FMLoRA 0.99
$\rho$ for EFMLoRA 0.1
$\beta$ for EFMLoRA 0.99
@@ B.2 Details on models
$\rho$ for FMLoRA 0.6
$\beta$ for FMLoRA 0.99
$\rho$ for EFMLoRA 0.1, 0.2, 0.5
$\beta$ for EFMLoRA 0.99
@@
Algorithm 1 Pseudocode of the FMLoRA
@@ B.3 Details on hyperparameters
$\rho$ for EFMLoRA 0.2,0.6,1
$\beta$ for EFMLoRA 0.99
@@
Figure 4: Evolution of balancedness across layers during training with Adam and FMLoRA.
@@ C. Algorithm
The two algorithms presented in 1 and 2 describe the training procedures of the proposed FMLoRA and its accelerated variant EFMLoRA.
Algorithm 2 Pseudocode of the EFMLoRA
@@ D.1 The approximate ability of EMA perturbation
D.2 The change in balancedness during FMLoRA training
We consider few shot learning with LoRA on RoBERTa-large. For dataset MNLI, 1st, 12th and 24th query layers’ $2|B_{t,l}|$ are plotted, where $t$ denotes the iteration and $l$ denotes the layer index. The layers are chosen to represent early, middle, and final stages of RoBERTa. Balancedness of FMLoRA and Adam on different layers are plotted in Fig. 4. Balancedness may increase or decrease across different layers. As shown in Fig. 4, the balancedness of FMLoRA in the first query layer of RoBERTa-large gradually decreases during training, while in the 12th layer, it first decreases and then increases. In contrast, the balancedness in the 24th layer continuously increases. An increase typically occurs when parameter magnitudes in both low-rank subspaces grow simultaneously. This behavior can be influenced by factors such as the learning rate, optimization algorithm, weight decay, and other regularization strategies. Despite these occasional increases, FMLoRA generally maintains lower balancedness than Adam in most layers, suggesting its capacity to induce implicit regularization during training.